# Enhanced Expressivity in Graph Neural Networks with Lanczos-Based Linear Constraints

### v1

@@ Introduction
The function $\bm{f}\colon S\cup\delta S\rightarrow\mathbb{R}$ represents the Neumann eigenvector satisfying the Neumann conditions. The vertex boundary, $\delta S$, of an induced subgraph consists of all vertices not in $S$ but adjacent to at least one vertex in $S$. Specifically, the first constraint encodes the link representation ($\sum_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y}\in S,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y}\sim{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}x}}(({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{f}(x)})-({\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\bm{f}(y)})=0$, $[{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}0},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}0},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1},{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}-1},{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}-1},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}]$ ; building on previous link prediction research, we consider nodes that are two hops away from the query nodes, where $S$ represents the one-hop-away nodes, and $\delta S$ denotes the boundary nodes between one-hop and two-hop-away nodes). This can be reformulated as:
$\min_{\bm{f}\in\mathbb{R}^{n}}\bm{f}^{\top}\mathbf{L}\bm{f},\text{ subject to }\lVert\bm{f}\rVert=1\text{ and }\mathbf{C}^{\top}\bm{f}=0,$ (4)
where $\mathbf{L}\in\mathbb{R}^{n\times n}$ is a symmetric and large sparse matrix, and $\mathbf{C}\in\mathbb{R}^{n\times l}$ (with $n\gg l$) is also large, sparse, and of full column rank. The time complexity involved in extracting subgraphs depends on the product of the maximum degree of nodes and the count of nodes within nodes in the boundary. When we enforce that the eigenvectors satisfy the constraints related to induced subgraphs and link representation, we ensure that the corresponding features adhere to these constraints.
@@
Constraints $\mathbf{C}$
The second subgraph extraction policy we propose is based on vertex-deleted subgraphs. The sufficient conditions under which LLwLC can solve graph isomorphism entails that LLwLC is a universal approximator of functions defined on graphs [14]. Given that we can encode any subgraph into our eigenbasis, we can examine whether a specific substructure collection can completely characterize each graph. By the reconstruction conjecture [45], we assume we can reconstruct the graph if we have all the $n-1$ vertex-deleted subgraphs. According to [26], exponentially many graphs are determined by their spectrum. Thus, it is believable that the same conjecture holds for the eigenbasis of graph Laplacian matrices of exponentially many graphs.
In extracting vertex-deleted subgraphs from a dense graph, the worst-case time complexity is $\mathcal{O}(n^{3})$. This complexity is due to the high number of edges in dense graphs, leading to increased computational load. For sparse graphs with fewer edges, the complexity is reduced by a factor of $\mathcal{O}(n)$. Our empirical results show that applying a few constraints significantly improves performance, making the practical overhead align linearly with the number of nodes. This approach is similar to the marker-based method [38], as both consider slightly perturbed input graphs.
@@
Furthermore, we conjecture that if the reconstruction conjecture holds [45] and the assumption that almost all graphs are reconstructable by their spectrum [26] is satisfied and the substructure collection contains all graphs of size $k=n-1$ with the form of ${\sum}_{x\in S}{}\bm{f}(x)d_{x}=0$ (for every vertex-deleted subgraph $S$), then it is believable and we conjecture that LLwLCNet can distinguish almost all non-isomorphic graphs of size $n$ and is therefore universal.
@@
In our study, we empirically discovered that effective outcomes can be attained with a limited number of constraints, such as Neumann constraints, or in another ablation study involving only ten vertex-deleted subgraphs. This observation aligns with the findings of Bollobás [10], who demonstrated that almost all graphs can be reconstructed using only three vertex-deleted subgraphs. Given the prohibitive cost of considering all $n$ vertex-deleted subgraphs, our approach, inspired by  [6], involves selecting a subset of constraints for consideration. During each epoch, we stochastically choose $k$ vertex-deleted subgraphs for analysis.
Experiments
To demonstrate the effectiveness of our method in addressing node automorphism and leveraging subgraphs, we conducted experiments focused on the link prediction task. We compared the performance of LLwLCNet against traditional heuristics (CN [5], RA [57], AA [1]), vanilla GNNs (GCN [25], SAGE [21]), GNNs modifying the input graph of MPNNs (SEAL [53], NBFNet [58]), and GNNs with manual features as pairwise representations (Neo-GNN [51], BUDDY [13]). Baseline results are from Chamberlain et al. [13]. Our evaluation includes five link prediction benchmarks: Cora, Citeseer, Pubmed [49] (Planetoid datasets), OGBL-Collab, and OGBL-Vessel [24]. Dataset statistics are shown in Appendix. Baseline results for OGBL-Collab and OGBL-Vessel are from the OGB leaderboard.
@@
In our experiments, we used a learning rate of 0.001 for 20 training epochs. The model has two 32-channel MLP layers, each with ReLU non-linearities and dropout. We cap the eigenpairs at 10, padding missing pairs with zeros. Both PyTorch [40] and PyTorch Geometric [17] were used for implementation. The binary cross entropy loss measured training loss. Following SEAL, we altered 10% of links for test data and used the remaining 90% for training.
@@
As depicted in Table 1, LLwLC stands out as a robust framework for link prediction, consistently delivering impressive performance on link prediction benchmarks with metrics given in the first row. With only 0.02M, LLwLC outperforms previous models on the Planetoid dataset, highlighting the significance of encoding subgraph structures and selective subsets of node relations for superior link prediction results. Our model also achieves state-of-the-art performance on the OGBL-Collab dataset with just 0.02M, compared to BUDDY (1.10M) and SEAL (0.50M). Increasing blocks to 0.03M yields a 67.50% HR@50 score. On OGBL-Vessel, we achieve competitive results with just 0.019M and 10% of the training data, demonstrating LLwLC’s effectiveness in a lightweight architecture.
@@
Table 2: Time Complexity Comparison
Time Complexity SEAL BUDDY NBFNet LLwLC
(b) Expressivity Models’ Time Complexity.
Time Complexity $k$-IGN $k$-GNN LLwLC GSN (3-WL)
Preprocessing - - $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{k})$
Sparse $\mathcal{O}(n^{k})$ $\mathcal{O}(n^{k})$ $\mathcal{O}(n)$ $\mathcal{O}(n)$
Time Complexity
The time complexity of our method is $\mathcal{O}(\kappa E)+\mathcal{O}(k^{2}n)+\mathcal{O}(k^{3})$ for the outer loop (Lanczos algorithm), the QR factorization, and the pseudo-inverse of $k\times k$ matrix, respectively, where $\kappa\ll n$ is the number of computed eigenvectors and $k\ll n$ is the number of linear constraints. Table 2(a) compares the time complexities of LLwLCNet and other link prediction methods.
Comparing the LLwLCNet model’s time complexity with $k$-WL expressive models provides valuable insights. It is conjectured that LLwLCNet is a universal approximator with a worst-case time complexity of $\mathcal{O}(n^{3})$ (where $k=n-1$). Aligned with [5] research and supported by our empirical experiments, it has been established that graph reconstruction can be effectively achieved with just a few constraints. This leads to time complexity linear to the number of nodes. Contrastingly, established $k$-WL expressive models, such as $k$-IGNs [30] and $k$-GNNs [34], are known to have a higher time complexity of $\mathcal{O}(n^{k})$.
Furthermore, we compare LLwLCNet’s time complexity with GSN [11]. It explores substructures by counting specific substructures. However, it relies on prior knowledge and inductive bias for task-specific substructure selection and deals with the subgraph isomorphism. In GSN, the preprocessing step, in general, is $\mathcal{O}(n^{k})$ for a generic substructure of size $k$. In contrast, our experimental results demonstrate that LLwLCNet can capture graph properties without requiring inductive bias. Comparisons are in Table 2(b).
@@
To demonstrate the impact of the applied constraints, we compare the eigenbasis estimation using the ground truth Laplacian matrix $\mathbf{L}$ (with edges provided) against the approximation of the eigenbasis using the ground truth Laplacian matrix $\mathbf{L}$ alongside Neumann constraints $\mathbf{C}$. This comparison is conducted on three benchmark datasets, as detailed in Table 3(a). We notice that including Neumann constraints, $\mathbf{C}$, markedly enhances the results.
In our extended research, we explore the impact of adding additional constraints to our framework. As detailed in Table 3(b), we explore the influence of integrating an increased number of constraints using vertex-deleted subgraphs. Our observations reveal that implementing merely ten constraints from vertex-deleted subgraphs yields state-of-the-art enhancements in benchmark datasets. Notably, within the OGBL-Collab dataset, the constraints significantly boost performance metrics (e.g., HR@50 increases from $42.83$ in LanczosNet, which does not apply linear constraints to the input graph, to $69.40$ with the incorporation of ten vertex-deleted subgraph constraints). Similar improvements are evident in the Cora dataset, with HR@100 rising from $90.80$ to $93.10$, and PubMed shows similar improved results over the baseline.Our results are consistent with the theoretical framework. It demonstrates that incorporating $n-1$ vertex-deleted subgraphs as linear constraints leads to a universal approximator function. Significantly, it aligns with [10], suggesting nearly all graphs can be reconstructed using three vertex-deleted subgraphs.
@@
Wall-time and Reduction Dataset
Table 4: Wall-time comparisons with LP models. Training time is one epoch.
@@
Related Work
@@
The expressivity of GNNs is typically expressed in terms of their ability to distinguish non-isomorphic graphs. As no polynomial-time algorithm for solving the graph isomorphism problem is known, developing GNNs that are both expressive and efficient poses a major challenge. Xu et al. [48] found that the expressivity of MPNNs is limited to that of the 1-WL test. This limitation is crucial in real-world applications, as the 1-WL test cannot distinguish certain graph structures, such as isomorphism in attributed regular graphs, measuring node distances, and counting cycles [28]. Recent studies have addressed these limitations with four approaches: adding random attributes to nodes [41], using deterministic positional features [53], developing higher-order GNNs to surpass the 1-WL test’s expressivity limits [31], and subgraph GNNs applying markings, such as the node-deletion approach of ESAN [6]. Please refer to the recent survey [33] for a comprehensive overview of these techniques. Our approach aligns with the direction of subgraph GNNs.
@@
Conclusion
In this work, we introduced the Learnable Lanczos algorithm with Linear Constraints (LLwLC), a novel method designed to enhance the expressivity of Graph Neural Networks (GNNs). Through the incorporation of two novel subgraph extraction strategies, we managed to construct a lightweight architecture that minimizes reliance on extensive training datasets. Empirical results show that our method significantly improves performance in link prediction tasks across various benchmark datasets. Notably, the LLwLC achieved 20$\times$ and 10$\times$ speedup, requiring only 5% and 10% data from the PubMed and OGBL-Vessel datasets respectively, compared to the state-of-the-art methods. These findings underscore the practical utility and theoretical advancement of our method, illustrating the LLwLC’s potential as a universal approximator and its ability to differentiate between $k$-regular graphs. The advancements made with the LLwLC not only represent a significant contribution to the field but also set a promising direction for future exploration and development in the realm of GNNs. As a further future work, we believe investigating the impact of learning linear constraints between nodes and edges within the input graph and encoding them to the eigenbasis of the graph Laplacian matrix, can promise further advancements in this field.

### v2

@@ 3.4 Subgraph Extraction Policy
The function $\bm{f}\colon S\cup\delta S\rightarrow\mathbb{R}$ represents the Neumann eigenvector satisfying the Neumann conditions. The vertex boundary, $\delta S$, of an induced subgraph consists of all vertices not in $S$ but adjacent to at least one vertex in $S$. Specifically, the first constraint encodes the link representation. Building on previous link prediction research, we consider nodes that are two hops away from the query nodes, where $S$ represents the one-hop-away nodes, and $\delta S$ denotes the boundary nodes between one-hop and two-hop-away nodes. It ensures that the aggregate of eigenvector differences across nodes equates to zero, as expressed by $\sum(\bm{f}(x)-\bm{f}(y))=0$. The column corresponding to the subgraph is constructed based on the equation. The degrees of the nodes involved in the constraint are entered into the column to satisfy the equation. Specifically, the degrees of nodes two-hop-away are negated, and the degrees of nodes one-hop-away are included only if they are connected to nodes two hops away. Entries for uninvolved nodes are set to 00.
Example 1.
Consider the graph , where the query nodes are orange, the nodes in $S$ are blue, and the boundary nodes $\delta S$ are green. The boundary constraint $\sum_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y}\in S,{\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}x}\in\delta S,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y}\sim{\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}x}}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{f}(x)}-{\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}\bm{f}(y)})=0$ is realized by the column $[{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}0},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}0},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1},{\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}-1},{\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}-1},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}]^{\top}$ of the matrix $\mathbf{C}$.
For the induced subgraph case, we populate its corresponding column in $\mathbf{C}$ with the degrees of each node (and fill the remaining entries of the column with 00 for nodes not involved in the subgraph), ensuring that the derived eigenvectors fulfill the condition $\sum\bm{f}(x)d_{x}=0$ (detailed in Vertex-deleted Subgraphs). This condition ensures that the features learned using this eigenbasis reflect the imposed constraints.
Thus, the Neumann eigenvalue problem can be reformulated as Equation 1 where $\mathbf{L}\in\mathbb{R}^{n\times n}$ is a symmetric and large sparse matrix, and $\mathbf{C}\in\mathbb{R}^{n\times l}$ (with $n\gg l$) is also large, sparse, and of full column rank. The time complexity involved in extracting subgraphs depends on the product of the maximum degree of nodes and the count of nodes in the boundary. When we enforce that the eigenvectors satisfy the constraints related to induced subgraphs and link representation, we ensure that the corresponding features adhere to these constraints.
@@
Vertex-deleted Subgraphs
The second subgraph extraction policy we propose is based on vertex-deleted subgraphs. Given that we can encode any subgraph into our eigenbasis, we can examine whether a specific substructure collection can completely characterize each graph. By the reconstruction conjecture [45], we assume we can reconstruct the graph if we have all the $n-1$ vertex-deleted subgraphs. According to Koval & Kwan [26], exponentially many graphs are determined by their spectrum. Thus, we propose that each column of the constraint matrix corresponds to a vertex-deleted subgraph, and we include the degrees of the nodes within the corresponding subgraph in each column, representing deleted nodes by zero entries.
Example 2.
Consider the graph , where the red nodes are deleted, and the orange nodes are contained in the subgraph. The constraint ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\sum\bm{f}(x)d_{x}=0}$ is represented by the column $[{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}2},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}2},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}2},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}2},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}]^{\top}$ in the matrix $\mathbf{C}$.
We build the other columns of the constraint matrix accordingly. The intuition behind defining such a constraint is that it implicitly represents the vertex deleted subgraph. Learning the filters over the eigenvalue matrix of the graph Laplacian leads to defining an affinity matrix based on Gaussian functions and driving the corresponding adjacency matrix based on the threshold [29]. Thus, $\mathbf{L}=\mathbf{I}-\mathbf{D}^{-1}\mathbf{W}$ where $\mathbf{W}$ is the affinity matrix [15]. $\mathbf{P}=\mathbf{D}^{-1}\mathbf{W}$ is the Markov transition matrix and its powers $\mathbf{P}^{t}$ are the probability of $t$-hop-away distance between the nodes $i$ to node $j$. Thus, the diffusion map is defined based on the right eigenvector and eigenvalues of the Markov transition matrix $\mathbf{P}$. If $\lambda_{i}$ and $\bm{\psi}_{i}$ are the eigenpairs of the matrix $\mathbf{P}$ then the eigenbasis of diffusion map after $t$-steps for the $i$th value is $\mathbf{\Phi}_{t}(i)=(\lambda_{1}^{t}\psi_{1}(t),\lambda_{2}^{t}\psi_{2}(t),...)$. Projecting to null space of the constraint matrix, $(\mathbf{I}-(\mathbf{C}^{\top}(\mathbf{C}\mathbf{C}^{\top})^{-1}\mathbf{C}))$ cancels out the nodes with zero entries and considers the others based on the degree we put in the constraint matrix. The diffusion map embedding adapts to reflect diffusion patterns that are aware of the vertex deletions.
In extracting vertex-deleted subgraphs from a dense graph, the worst-case time complexity is $\mathcal{O}(n^{3})$ due to the high number of edges in dense graphs, leading to increased computational load. For sparse graphs with fewer edges, the complexity is reduced by a factor of $\mathcal{O}(n)$. Our empirical results show that applying a few constraints significantly improves performance, making the practical overhead align linearly with the number of nodes. This approach is similar to the marker-based method [38], as both consider slightly perturbed input graphs.
@@
In our study, we empirically discovered that effective outcomes can be attained with a limited number of constraints, such as Neumann constraints, or in another ablation study involving only ten vertex-deleted subgraphs. This observation aligns with the findings of Bollobás [10], who demonstrated that almost all graphs can be reconstructed using only three vertex-deleted subgraphs. Given the prohibitive cost of considering all $n$ vertex-deleted subgraphs, our approach, inspired by [6], involves selecting a subset of constraints for consideration. At each epoch, we stochastically choose $k$ vertex-deleted subgraphs for analysis.
Table 1: Time Complexity Comparison
Time Complexity SEAL BUDDY NBFNet LLwLC
(b) Expressivity Models’ Time Complexity.
Time Complexity $k$-IGN $k$-GNN LLwLC GSN (3-WL)
Preprocessing - - $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{k})$
Sparse $\mathcal{O}(n^{k})$ $\mathcal{O}(n^{k})$ $\mathcal{O}(n)$ $\mathcal{O}(n)$
Time Complexity
The time complexity of our method is $\mathcal{O}(\kappa E+k^{2}n+k^{3})$ for the outer loop (Lanczos algorithm), the QR factorization, and computing the pseudo-inverse of $k\times k$ matrix, respectively, where $\kappa\ll n$ is the number of computed eigenvectors and $k\ll n$ is the number of linear constraints. Table 1(a) compares the time complexities of LLwLCNet and other link prediction methods.
Aligned with [5] research and supported by our empirical experiments, it has been established that graph reconstruction can be effectively achieved with just a few constraints. This leads to time complexity linear to the number of nodes. Contrastingly, established $k$-WL expressive models, such as $k$-IGNs [30] and $k$-GNNs [33], are known to have a higher time complexity of $\mathcal{O}(n^{k})$.
Furthermore, we compare LLwLCNet’s time complexity with GSN [11], which improves expressivity by counting specific substructures. However, it relies on task-specific substructure selection to introduce a suitable inductive bias and deals with the subgraph isomorphism. In GSN, the preprocessing step, in general, is $\mathcal{O}(n^{k})$ for a generic substructure of size $k$. In contrast, our experimental results demonstrate that LLwLCNet can capture graph properties without requiring task-specific prior knowledge. Comparisons are in Table 1(b).
4 Experiments
To demonstrate the effectiveness of our method in addressing node automorphism and leveraging subgraphs, we conducted experiments focused on the link prediction task. We compared the performance of LLwLCNet against traditional heuristics (CN [5], RA [57], AA [1]), vanilla GNNs (GCN [25], SAGE [21]), GNNs modifying the input graph of MPNNs (SEAL [53], NBFNet [58]), and GNNs with manual features as pairwise representations (Neo-GNN [51], BUDDY [13]). Baseline results are from Chamberlain et al. [13]. Our evaluation includes five link prediction benchmarks: Cora, Citeseer, Pubmed [49] (Planetoid datasets), OGBL-Collab, and OGBL-Vessel [24]. Dataset statistics are shown in the appendix. Baseline results for OGBL-Collab and OGBL-Vessel are from the OGB leaderboard.
@@ 4 Experiments
In our experiments, we used a learning rate of 0.001 for 20 training epochs. The model has two 32-channel MLP layers, each with ReLU non-linearities and dropout. We cap the eigenpairs at 10, padding missing pairs with zeros. Both PyTorch [40] and PyTorch Geometric [17] were used in our implementation. The model was trained using the binary cross entropy loss. Following SEAL, we altered 10% of links for test data and used the remaining 90% for training.
@@
As depicted in Table 2, LLwLC stands out as a robust framework for link prediction, consistently delivering high performance on link prediction benchmarks with metrics given in the first row. With only 0.02M parameters, LLwLC outperforms previous models on the Planetoid dataset, highlighting the significance of encoding subgraph structures and selective subsets of node relations for superior link prediction results. Our model also achieves state-of-the-art performance on the OGBL-Collab dataset with just 0.02M, compared to BUDDY (1.10M) and SEAL (0.50M). Increasing blocks to 0.03M yields a 67.50% HR@50 score. On OGBL-Vessel, we achieve competitive results with just 0.019M and 10% of the training data, demonstrating LLwLC’s effectiveness in a lightweight architecture.
@@
To demonstrate the impact of the applied constraints, we compare the eigenbasis estimation using the ground truth Laplacian matrix $\mathbf{L}$ (with edges provided) against the approximation of the eigenbasis using $\mathbf{L}$ alongside Neumann constraints $\mathbf{C}$. This comparison is conducted on three benchmark datasets, as detailed in Table 3(a). We notice that including Neumann constraints, $\mathbf{C}$, markedly enhances the results.
In our extended research, we explore the impact of adding additional constraints to our framework. As detailed in Table 3(b), we explore the influence of integrating an increased number of constraints using vertex-deleted subgraphs. Our observations reveal that implementing merely ten constraints from vertex-deleted subgraphs yields state-of-the-art enhancements in benchmark datasets. Notably, within the OGBL-Collab dataset, the constraints significantly boost performance metrics (e.g., HR@50 increases from $42.83$ in LanczosNet, which does not apply linear constraints to the input graph, to $69.40$ with the incorporation of ten vertex-deleted subgraph constraints). Similar improvements are evident in the Cora dataset, with HR@100 rising from $90.80$ to $93.10$, and PubMed shows similar improved results over the baseline. Our results are consistent with the theoretical framework. It significantly aligns with [10], suggesting nearly all graphs can be reconstructed using three vertex-deleted subgraphs.
@@
Table 4: Wall-time comparisons with LP models. Training time is measured for one epoch. The LLwLC model demonstrates state-of-the-art performance in both benchmark datasets while reducing the training data (Reduction) requirement by up to $90\%$, using only $10\%$ of the data.
@@
Wall-time and Reduction Dataset
5 Related Work
@@
The expressivity of GNNs is typically expressed in terms of their ability to distinguish non-isomorphic graphs. As no polynomial-time algorithm for solving the graph isomorphism problem is known, developing GNNs that are both expressive and efficient poses a major challenge. Xu et al. [48] found that the expressivity of MPNNs is limited to that of the 1-WL test. This limitation is crucial in real-world applications, as the 1-WL test cannot distinguish certain structures, such as regular graphs, and does not capture several natural graph properties well, such as distances and cycle counts [28]. Recent studies have addressed these limitations with four approaches: adding random attributes to nodes [41], using deterministic positional features [53], developing higher-order GNNs to surpass the 1-WL test’s expressivity limits [31], and subgraph GNNs applying markings, such as the node-deletion approach of ESAN [6]. Please refer to the recent survey by Morris et al. [34] for a comprehensive overview of these techniques. Our approach aligns with the direction of subgraph GNNs.
@@
6 Conclusion
In this work, we introduced the Learnable Lanczos algorithm with Linear Constraints (LLwLC), a novel method designed to enhance the expressivity of Graph Neural Networks (GNNs). Through the incorporation of two novel subgraph extraction strategies, we managed to construct a lightweight architecture that minimizes reliance on extensive training datasets. Empirical results show that our method significantly improves performance in link prediction tasks across various benchmark datasets. Notably, the LLwLC achieved 20$\times$ and 10$\times$ speedup, requiring only 5% and 10% data from the PubMed and OGBL-Vessel datasets respectively, compared to the state-of-the-art methods. These findings underscore the practical utility and theoretical advancement of our method, illustrating the LLwLC’s potential as a more expressive approach than 2-WL and its ability to differentiate between $k$-regular graphs. The advancements made with the LLwLC not only represent a significant contribution to the field but also set a promising direction for future exploration and development in the realm of GNNs. As a further future work, we believe investigating the impact of learning linear constraints between nodes and edges within the input graph and encoding them to the eigenbasis of the graph Laplacian matrix, can promise further advancements in this field.